#1. In a class, there are 20 men and 15 women. Find the ratio of the number of men to the number of students in the class. Express the ratio as a fraction reduced to the lowest term.

2. Determine whether each of the ordered pair is a solution of,
3x-4y>7: (0,0), (3,-6), (-2,-5), -3,4)
#3 A chemist needs to mix a 40% acid solution with 70% acid solution to obtain the 12 litters of 50%acid solution. How many liters of each of the acid solutions must be used?
#4. Solve the proportion:45/2=360/x

Answer by stanbon(75887)

In a class, there are 20 men and 15 women. Find the ratio of the number of men to the number of students in the class. Express the ratio as a fraction reduced to the lowest term.
men:women = 20:15 = 4:3
----------------------------------
2. Determine whether each of the ordered pair is a solution of,
3x-4y>7: (0,0), (3,-6), (-2,-5), (-3,4)
(0,0): 0-0 > 7 false
(3,-6): 9-24 > 7 false
(-2,-5): -6 + 20 > 7 true
(-3,4): -9 -16 > 7 false
===============================

#3 A chemist needs to mix a 40% acid solution with 70% acid solution to obtain the 12 liters of 50% acid solution. How many liters of each of the acid solutions must be used?
-----
Equation:
acid + acid = acid
0.40x + 0.70(12-x) = 0.50*12
Multiply thru by 100 to get:
40x + 70*12 -70x = 50*12
-30x = -20*12
x = 8 liters (amt. of 40% solution in the mixture)
12-8 = 4 liters (amt. of 70% solution in the mixture)
===========================================================
#4. Solve the proportion:45/2=360/x
Cross-multiply
45x = 2*360
x = 16
=============
Cheers,
Stan H.